The police use the formula v=4.9 square root L to estimate the speed of a car v based on the length L of its skid marks on dry pavement.
a. What is a reasonable estimate for a car's speed if its skid marks measure 220 feet?
b. What would be the length of your skid marks if you drove with the speed of 65mph?
Can someone explain this to me?

Answer by checkley77(12844)

V=4.9*SQRTL
V=4.9*SQRT200
V=4.9*14.14
V=69.29 MPH. ANS.
65=4.9*SQRTL
65/4.9=SQRTL
13.2653=SQRTL
L=13.2653^2
L=175.97 FEET IS THE LENGTH OF THE SKID MARKS.